OpenStudy (anonymous):

find the value of the x in the expression (2a^4b^2)^x=4a^8b^4

OpenStudy (anonymous):

\[\huge (2a^4.b^2)^x = 4a^8.b^4\] \[\huge (2a^4.b^2)^{x} = (2a^4.b^2)^2\] \[\huge x^a = x^b \implies a = b\] So, \[\huge \color{green}{x = 2}\]

OpenStudy (anonymous):

thanks for the help